OpenStudy (anonymous):

if 5 out of 25 people like chocolate ice cream what percent of the poeple do NOT like chocolate ice cream choices first to give me the answer i will fan give best response 20% 5% 80% 30% plz just tell me the answer

OpenStudy (anonymous):

opps 5 out of 25

OpenStudy (tkhunny):

Please just read the code of conduct and heed its words. The second one should be 50%?

OpenStudy (tkhunny):

Lacking those 5, how many are left?

OpenStudy (anonymous):

okay i fixed it

OpenStudy (anonymous):

5 out of 25

OpenStudy (tkhunny):

5 of 25 like. What number of 25 don't like?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

No you need to FIND the people that dont like chocolate ice cream

OpenStudy (tkhunny):

We don't need to write down their names. Just count them. How many are there?

OpenStudy (anonymous):

15 people dont like chocolate ice cream

OpenStudy (tkhunny):

5 + 15 = 25? Are you sure?

OpenStudy (anonymous):

10

OpenStudy (tkhunny):

5 + 10 = 25? Are you sure?

OpenStudy (anonymous):

20

OpenStudy (tkhunny):

5 + 20 = 25? Are you sure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am in a rush plz i just need the answer

OpenStudy (tkhunny):

I don't actually care if you are in a hurry. I'm interested in you learning. If you don't have the time, feel free to come back when you do. 5 like 20 don't like 25 is the total 5 is what percent of 25? This is the percentage that likes choc ice cream. 20 is what percent of 25? This is the percentage that does not like choc ice cream.

OpenStudy (anonymous):

5 is 20 percent of 25 20 is what percent of 25? it is 5 @tkhunny

OpenStudy (tkhunny):

Oi, oi, oi. I think you are just guessing. How did you get 5 is 20% of 25? Did you just pull that out of a hat or did you calculate it, somehow?

OpenStudy (anonymous):

oi oi oi claculator

OpenStudy (tkhunny):

This is a great puzzle. Calculator: 5/25 = 0.20 = 20% What Method: 20/25 = 5%?? Please try to be consistent and reasonable.

OpenStudy (anonymous):

regardless i know the answer and the work

OpenStudy (anonymous):

5/25=.2 So 20% of the people like it. Total is 100%, so 100-20=80% :DDD @tkhunny

OpenStudy (tkhunny):

Okay, so where did that "5" come from , yesterday? It's not a matter of "regardless" or "know the answer". It's a matter of being consistent and consistently correct. If you EVER get to a point where you just start throwing out numbers, maybe you need some sleep, or something. Good work on the right answer.